Kieth has 83 coins in nickels and dimes. He has a total of $6.95. How many dimes and how many nickels does Keith have
Answer by addingup(3677)

n+d= 83 subtract d both sides:
n = 83-d we'll use this value for n next:
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
0.05n+0.10d = 6.95
0.05(83-d)+0.10d = 6.95
4.15-0.05d+0.10d = 6.95 Subtract d on left and 4.15 on both sides:
0.05d = 2.80 divide both sides by 0.05
d = 56 he has 56 dimes and
83-56 = 27 nickels
``````````````````````````````````````````````
Let's check:
56 dimes = 5.60
27 nickels= 1.35
Total...... 6.95 We have the correct answer.
J